OpenStudy (anikate):

find the sum of the first 20 terms: 1+ 1/2+ 1/4+ 1/8+... Check my answer I got: 1

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anonymous):

Here's the formula. \[S_n = \frac{a_1(1 - r^n)}{1 - r}\]

OpenStudy (anonymous):

Check it once more.

OpenStudy (anonymous):

\[\large S_n = \frac{(1)(1 - (\frac{1}{2})^{20})}{1 - \frac{1}{2}}\]

OpenStudy (anonymous):

\[S_{20} = \frac{0.999}{0.5}\]

OpenStudy (anikate):

thx guys, I didnt reply early enough because my laptop died